OpenStudy (anonymous):

The average of 8 whole numbers is 8. If 7 of these numbers are 1, the the eighth number is ______.

Directrix (directrix):

If 8 numbers average to 8, then the sum of the numbers is 64. 64 - the 7 ones = 57 for the eighth number.

OpenStudy (radar):

(7+x)/8=8 the way an average is obtained. 7+x = 64 multiplying thru by 8 x=64-7=57 Just as direcrix has described.